# Portable Targeted Sampling Framework Using LLVM

### v1

@@ IV-B2 Sample Validation
Figure 5 presents the prediction errors of nuggets generated by the Random and K-means approaches for each workload on the real-hardware machines listed in Table I. As shown, many workloads exhibit significant errors, often exceeding the accuracy reported by prior methodologies such as SimPoint (3% for single-threaded workloads) and LoopPoint (2.9% for 8-threaded workloads).
Several factors contribute to these elevated errors. (1) The sampling methodologies are naively designed. For example, the Random approach relies heavily on luck and the specific composition of the interval pool. (2) Previous work did not validated samples on real hardware. Instead, they rely on user-mode-only simulation, which is typically less noisy due to more controllable thread interleaving compared to actual hardware. In Section V-A, we further analyze the sources of prediction error using smaller workloads in simulation. (3) Prior work did not validate large workloads as we used in our experiments. Larger and more realistic applications exhibit increasingly complex behaviors, which demand more representative sample sets, as discussed in Section II-C. In such cases, large program phases can amplify prediction error. (4) Hardware-based nuggets incur additional overhead for the begin and end interval markers, particularly in multi-threaded contexts, which further inflates prediction error. We provide a detailed discussion of this overhead in Section IV-B3.
@@ IV-B3 Hook Overhead
Nugget uses hook instructions to mark sample intervals, which can introduce runtime overhead during workload execution. In Figure 5, the “is” benchmark illustrates how validation results can be distorted by this hook overhead. Because “is” frequently executes only a small number of IRBBs within a parallel region, it triggers an excessive number of hook invocations. “is” was also one of the workloads with the larges overheads for analysis as shown in Figure 2. This leads to substantial overhead and results in inaccurate prediction errors.
@@ V-A1 Experimental Setup
TABLE III: gem5 Simulation Model Configuration
We use the selected nuggets to conduct experiments on both gem5, using full-system simulation, and real hardware. In addition to the systems listed in Table I, we include two additional real hardware platforms: an Intel i9-9900X system and an Intel i7-9700 system. The simulation models used are listed in Table III and run a guest Ubuntu 24.04 kernel, matching the software environment of the real hardware.
@@ V-A2 Results
To isolate the sources of prediction error, we compare three types of platform pairs: (1) systems with different ISAs but small micorarchiture differences, (2) systems with different significantly different microarchitectural configurations, but the same ISA, and (3) systems that differ in both ISA and microarchitecture.
@@ V-B Model Accuracy Evaluation
In this section, we present how we used nuggets to identify sources of errors in gem5’s models. Identifying sources of errors in simulation model is key to establishing baseline performance for evaluating architectural ideas.
@@
Figure 10 shows a comparison of the the number of memory instructions in our select nuggets for three different microcodings. The orange bars show the number of memory instructions with the default microcoding of gem5; the blue bars show the number of memory instructions if the paired memory instructions were microcoded as one microop; while the green bars show the number of memory instructions measured on native hardware. The blue bars closely track the green bars, suggesting that the microcoding of paired memory instructions in gem5 is different from the native hardware.
@@ VI Conclusion
In this paper we presented Nugget: a framework for building program snippets that can be used for evaluating architectural ideas using full-system simulators. Nugget enables automatic instrumentation of a workload’s binary for analysis, and for measurement. Binaries used for analysis can be run on real hardware much faster than prior work while the binaries used for measurement can be run on real hardware and simulator alike. Lastly, Nugget allows users to leverage real hardware to evaluate how well they represent the original workload.
Nugget is positioned to allow researchers to simulate workloads larger than the traditional sizes enabling researchers to gain insights that was not possible before.
@@ Acknowledgements
We thank Alen Sabu, Christopher Batten, Matthew D. Sinclair, Oscar Hernandez, and the Davis Computer Architecture Research Group for feedback used to improve this work. This work was supported by NSF grant 2311888, Department of Energy, Oak Ridge National Laboratory, and Los Alamos National Laboratory.

### v2

@@ IV-B2 Sample Validation
Figure 5 presents the prediction errors of nuggets generated by the Random and K-means approaches for each workload on the real-hardware machines listed in Table II. As shown, many workloads exhibit significant errors, often exceeding the accuracy reported by prior methodologies such as SimPoint (3% for single-threaded workloads) and LoopPoint (2.9% for 8-threaded workloads).
Several factors contribute to these elevated errors. (1) The sampling methodologies are naively designed. For example, the Random approach relies heavily on luck and the specific composition of the interval pool. (2) Previous work rarely validates samples on real hardware [26, 22, 41, 39, 49, 40, 36, 32, 14, 50, 30]. Instead, they mostly rely on user-mode-only simulation, which is typically less noisy due to more controllable thread interleaving compared to actual hardware [3, 9, 1, 17]. In Section V-A, we further analyze the sources of prediction error using smaller workloads in simulation. (3) Prior work rarely validates large workloads as we used in our experiments (e.g., reference input sizes for CPU2017). Larger and more realistic applications exhibit increasingly complex behaviors [51, 48, 41, 31, 26, 39], which demand more representative sample sets, as discussed in Section II-C. In such cases, large program phases can amplify prediction error. (4) Hardware-based nuggets incur additional overhead for the begin and end interval markers, particularly in multi-threaded contexts, which further inflates prediction error. We provide a detailed discussion of this overhead in Section IV-B3.
@@ IV-B3 Hook Overhead
Figure 6: The number of marker hook executions for the randomly selected nuggets shown in Figure 5. The number of hook executions is normalized to the total number of IRBB executions during all nugget executions for each workload and input pair.
Nugget uses hook instructions to mark sample intervals, which can introduce runtime overhead during workload execution. In Figure 6, we present the number of times the marker hooks are executed for the randomly selected nuggets shown in Figure 5. The number of hook executions is normalized to the total number of IRBB executions during all nugget executions for each workload and input pair. As shown, “is” exhibits a significantly higher number of hook executions compared to other workloads, indicating that the prediction error observed for “is” in Figure 5 is influenced by the excessive overhead from hook executions. Because “is” frequently executes only a small number of IRBBs within a parallel region, it triggers an excessive number of hook invocations. “is” was also one of the workloads with the largest overheads for analysis as shown in Figure 2. This leads to substantial overhead and results in inaccurate prediction errors. Hook overhead can only slow execution, so it can cause false negatives (rejecting good samples) but is unlikely to cause false-positive conclusions.
For other workloads, the number of hook executions does not show a clear correlation with the prediction error observed in Figure 5. Cleanly separating the impact of hook overhead from other sources of prediction error is still challenging. The number of hook executions and the prediction errors observed suggest that the cutoff line for a good marker should be executing lower than 10% for single-threaded workloads and lower than 1% for multithreaded workloads, in terms of the total IRBBs executed in the nugget, to avoid significant overhead impacting the validation results. Multithreaded workloads are more sensitive to hook overhead due to the synchronization required for counting and threshold checks, as discussed in Section III-D2.
@@ V-A1 Experimental Setup
TABLE III: gem5 Simulation Model Configuration. The large, small, and base out-of-order cores are internal models that target different performance characteristics.
We use the selected nuggets to conduct experiments on both gem5, using full-system simulation, and real hardware. In addition to the systems listed in Table II, we include two additional real hardware platforms: an Intel i9-9900X system and an Intel i7-9700 system. The simulation models used are listed in Table III and run a guest Ubuntu 24.04 kernel, matching the software environment of the real hardware.
@@ V-A2 Results
To isolate the sources of prediction error, we compare three types of platform pairs: (1) systems with different ISAs but small microarchitecture differences, (2) systems with different significantly different microarchitectural configurations, but the same ISA, and (3) systems that differ in both ISA and microarchitecture.
@@ V-B Model Accuracy Evaluation
In this section, we present how we used nuggets to identify sources of errors in gem5’s models. Identifying sources of errors in the simulation model is key to establishing baseline performance for evaluating architectural ideas.
@@
Figure 11 shows a comparison of the number of memory instructions in our selected nuggets for three different microcodings. The orange bars show the number of memory instructions with the default microcoding of gem5; the blue bars show the number of memory instructions if the paired memory instructions were microcoded as one microop; while the green bars show the number of memory instructions measured on native hardware. The blue bars closely track the green bars, suggesting that the microcoding of paired memory instructions in gem5 is different from the native hardware.
@@ VI Conclusion
In this paper we presented Nugget: a framework for building program snippets that can be used for evaluating architectural ideas using full-system simulators. Nugget enables automatic instrumentation of a workload’s binary for analysis, and for measurement. Binaries used for analysis can be run on real hardware much faster than prior work while the binaries used for measurement can be run on real hardware and simulators alike. Lastly, Nugget allows users to leverage real hardware to evaluate how well the selected samples represent the original workload.
Nugget is positioned to allow researchers to simulate workloads larger than the traditional sizes, enabling researchers to gain insights that were not possible before.
@@ Acknowledgements
We thank Alen Sabu, Christopher Batten, Derrick Quinn, Matthew D. Sinclair, Oscar Hernandez, and the Davis Computer Architecture Research Group for feedback used to improve this work. This work was supported by NSF grant 2311888, Department of Energy, Oak Ridge National Laboratory, and Los Alamos National Laboratory.